UT Mar 06 2022 - Split Night PETS/PEPSI - DEC




Night Info

  • Sunset 18:20
  • Nautical Twilight Ends 19:13
  • Astronomical Twilight Ends 19:42
  • Moon sets 21:39
  • Handover to PEPSI/PETS 20:06
  • PETS-7 20:36-02:40
  • Handover from PEPSI/PETS 02:41
  • Astronomical Twilight Starts 05:19
  • Nautical Twilight Stars 05:48
  • Sunrise 06:41

Staff

  • OSA: J. Williams
  • ISp: E. Solheid
  • Mtn Manager: J. Protho
  • ISA/Observer: B.Rothberg, R.T.Gatto
  • AO: J.C.Guerra

Night Plan

  • Staring Config Bino-LUCIs - authorize LUCI-1 AO
  • OSCO until 19:42
  • Short AO test 19:43-20:06
  • Bin 3 IQ test
  • LUCI-1 AO/ESM Testing 02:42-twlight
  • /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/
    • AO OT Version 76
    • ESM OT Version 36
  • Bin 2 ESM w/new AO Supervisor
    • M67 - twilight-02hr
    • Dual BLR - 02hr-twilight
  • Close separation targets (0.1"-1")
    • INAF Target J0732+3533 - 2" separation 1.5mag difference - could test this script on-sky beforehand -its vsisble first part of night (could roll over to Sat night)
    • BD+33 2137 RA10-11hr Star 1 in Binary folder R=10.3 - separation 2.4" separation 0.5 mag flux difference (might be too far apart - but wiki log showed this was an issue - but seeing was unsteady)
    • RA12h-13hr Star 1 in Binary Folder BD+00 2888 R=10.7 - separation 0.225"
    • HIP 65069 - Ra13hr+17Dec R=6.9 0.12" separation - need 64x2048 sub-frame (use to also test sub-framing)
    • 2MASSJ1128+3136 R=13.5 - 1" separation - 1mag difference in flux
    • 2MASSJ13061537+2043444 R=8.5 1.6" separation (64x2048 sub-framing) - 1.2mag difference in flux
  • Bin 3 IQ
    • bright
    • faint
  • Check on Bin 2 Full AO IQ
  • Continue to test RTD Sub-frame
  • Verify no jumps in AO w/LUCI N30 - IT 8578/IT8579
    • Repeat Observing RA14hr-15hr Star 1 R=10.2 BD +00 3109 (saw jump on this before)
    • S255 (UVa target) - twilight-midnight
  • Finish on Cat's Eye Nebula (FeII and Pa-Beta if possible)

Poor Weather Backup

  • LUCI PhotStd - if clear and poor seeing (2" or greater - can convert to track mode)
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
  • Rothberg Mergers
    • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V41.xml
      • NGC 4004 - 21hr-twilight
      • NGC 3597 - 23hr-02hr
      • UGC 9829 - 01:30hr-twilight
      • UGC 10607 - 02hr -twilight
      • UGC 10675 - 02hr-twilight


Night Log

Closed Dome

  • UT 00:02 - starting up LUCI GUIs, init all and taking test darks (L1 0013-0014, L2 0001-0002)
  • UT 00:11 - checking N30Fieldstop - its close, so adjusting - L1 0015-0017
  • UT 01:03 - winds are at the 25m/s level - clouds have cleared out for now
  • UT 02:45 - winds still at 23-25m/s level - humidity is creeping back up again - now 72%
  • UT 03:00 - Reconfiguring for Bino-PEPSI
  • UT 04:20 - winds up to 26.5 m/s - humidity now at 85%
  • UT 04:51 - humidity 91%
  • UT 05:34 - stars vanishing, must be the space dragon...
  • UT 05:41 - enclosure is icing up.....
  • UT 05:58 - J.Williams calling the night....D.Miller will work on NCPA data collection in about an hour
  • UT 10:01 - stopping NCPA work - had connections to summit freeze for both B.Rothberg and D.Miller a few times. Better to be safe in case connection drops. Not sure if the reading of 75m/s is correct on the info page.
  • UT 10:02 - D.Miller reverting to stable AO supervisor

-- BarryRothberg - 28 Feb 2022
Topic revision: r6 - 06 Mar 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback